# The Stochastic Schwinger Effect

### v1

@@ 2.1 Vacuum Instability and Pair Production
where the D’Alembertian is $\Box\equiv\partial^{\mu}\partial_{\mu}$. The bosonic functional determinant arises from integrating out charged scalars, while the fermionic one corresponds to integrating out Dirac fields and includes an additional trace over spinor indices. A convenient tool to evaluate these is the proper-time representation (see ˜97). For the scalar case, one has
@@ 3 Stochastic Schwinger Effect
$\Gamma_{\text{1-loop}}^{b}[A_{\mu}]$ $=\text{Tr}\ln\left[\frac{(\hat{p}-gQ\hat{A}({x}))^{2}+m^{2}-i\epsilon}{\hat{p}^{2}+m^{2}-i\epsilon}\right].$ (26)
@@ 3.1 Stationary background
By inserting ˜117 into ˜31, the decay probability appearing in ˜31 takes the form
@@
where, for shorthand, we denote $\int_{q}\equiv\int d^{4}q$. Here we summarise only the final expressions for the decay into scalar particles; the detailed derivation leading to this result is presented in Appendix B. Using ˜134 and 135, we obtain the one-loop vacuum decay probability into charged bosons induced by a stationary, stochastic gauge field as
@@
where $q^{\mu}=(\omega,\bm{q})$. Using ˜137, the corresponding spectral number density for fermionic pair production takes the form
@@ 3.2 Non-stationary background
with $\langle\!\!\langle\hat{F}_{\mu\nu}(t^{\prime\prime},\bm{q})\,\hat{F}^{\mu\nu}(t^{\prime},-\bm{q})\rangle\!\!\rangle=(q^{2}\eta^{\mu\nu}-q^{\mu}q^{\nu})\langle\!\!\langle\hat{A}_{\mu}(t^{\prime\prime},\bm{q})\,\hat{A}_{\nu}(t^{\prime},-\bm{q})\rangle\!\!\rangle$. The fermionic case follows from ˜137 as
@@ 4.1 Electromagnetic modes in a cold medium
We now compute the vacuum decay rate into a scalar species of mass $m$ in the presence of a cold–plasma background. This quantity admits exact analytical solutions in terms of the Gaussian hypergeometric function $_{2}F_{1}(a,b;c;z)$. The full derivation and closed forms are presented in Appendix C (see ˜145). For phenomenological clarity, here we report only the simplified expression obtained in the representative regime $\Lambda_{\mathrm{IR}}/\omega_{p}\simeq 1$ and $\Lambda_{\mathrm{UV}}/\omega_{p}\gg 1$; the final result reads
@@ 4.2 Dark photon background
Dark photons are hypothetical massive spin-1 bosons from a hidden $U(1)^{\prime}$ sector, kinetically mixed with the Standard Model hypercharge (see Fabbrichesi_2021 ; Caputo:2021eaa for reviews). In astrophysical and cosmological contexts, stochastic dark-photon fields may form with sizable energy densities, acting as classical sources for charged-particle production and providing a natural setting for the stochastic Schwinger effect. A massive vector field of mass $m_{A^{\prime}}$ propagates three physical degrees of freedom: two transverse and one longitudinal, with dispersion relation
@@
We now evaluate the vacuum decay rate into a scalar species of mass $m$ in the presence of a cold–plasma background. This quantity also admits an exact analytical expression in terms of the Gaussian hypergeometric functions; the detailed derivation and the resulting closed-form expressions can be found in Appendix C (see ˜146).
@@
The charged particles coupled to the dark photon may belong either to a dark sector, consisting of fermionic or bosonic states, or to the Standard Model, in which case they carry a small effective (millicharge) coupling $gQ=\varepsilon e$. In the dark-sector case, the coupling $gQ$ can naturally be of order unity, allowing a stochastic dark-photon background with mass $m_{A^{\prime}}$ to produce pairs with masses up to $m=m_{A^{\prime}}/2$. For SM particles, however, current experimental and astrophysical constraints on $\varepsilon$ impose stringent limits on such interactions. Collider and electroweak precision measurements yield $\varepsilon\lesssim 10^{-2}\,\text{–}\,10^{-3}$ for $m_{A^{\prime}}$ in the GeV range, while at higher masses, $m_{A^{\prime}}\gtrsim 100~\mathrm{GeV}$, these bounds become less restrictive Fabbrichesi:2020wbt .
@@ 5 Stochastic Schwinger in Axion–Gauge Field Reheating
where the infrared cut-off corresponds to modes residing within the Hubble horizon. In this work, we focus on early-time dynamics with $\mu_{\bm{q}}t\ll 1$, a regime that captures the most relevant phenomenology while avoiding backreaction. Within the Gaussian-window STFT framework in (41), this corresponds to imposing $\sigma\lesssim\mu_{\bm{q}}^{-1}$. As discussed in section˜3.2, in this context the window width is naturally momentum-dependent, scaling with the oscillation timescale of each mode, $\sigma_{\bm{q}}\sim\omega_{\bm{q}}^{-1}$, in close analogy with adaptive wavelet methods. From ˜40 and 41, together with ˜101 and the identification $\sigma\simeq\mu_{\bm{q}}^{-1}$, the gauge mode functions simplify to
@@
As we are working in the small time domain, we can approximate the periodic function as being nearly constant, $c(t,\bm{q})\approx 1$. Using the above, ˜135, ˜136, and in the early-time regime, i.e. $\xi\lvert\bm{q}\rvert t\ll 1$, we find
@@
where $y=\omega/|\mathbf{q}|$, with $e\simeq 2.7$ and the maximal momentum $q_{\text{max}}=m_{\chi}\xi$ set by the instability window in ˜82. The above expression does not admit a closed-form analytic solution. However, we were able to approximate its behaviour using complementary error functions as detailed in appendix˜C of this work. The explicit form of the spectral number density is presented in ˜155. Its contribution remains negligible for small values of $|\bm{q}|$ and becomes relevant for $|{\bm{q}}|\gtrsim 2m/\xi$, where it takes the following form
@@
By performing the momentum integral, we obtain the number density given in ˜156. In the regime of interest, $\xi\gtrsim 1$, it becomes
@@
Finally, we find the background decay rate to bosons as (see ˜158)
@@ Appendix A Mathematical tools
These identities hold for generic parameters by analytic continuation, except at gamma-function poles. For n=2n=-2, the integral in ˜106 evaluates to
@@ Appendix B Technical Details of Stochastic Effective Action
$\begin{split}1=&\exp\left[\text{Tr}\ln\left(\hat{\mathbb{I}}-\frac{gQ\,\hat{f}(A)}{\hat{p}^{2}+m^{2}-i\epsilon\hat{p}^{0}}\right)\right]\\ =&\exp\left[\text{Tr}\ln\left(\hat{\mathbb{I}}-\frac{gQ\,\hat{f}(A)}{\hat{p}^{2}+m^{2}-i\epsilon}+igQ\hat{f}(A)\,\hat{\delta}^{-}\right)\right]\\ =&\exp\left[\text{Tr}\ln\left(\left(\hat{\mathbb{I}}-\frac{gQ\,\hat{f}(A)}{\hat{p}^{2}+m^{2}-i\epsilon}\right)\,(1+i\hat{O}(A)\,\hat{\delta}^{-})\right)\right]\end{split}$ (119)
@@
We now turn to the proof of ˜30. Starting from
[…]
we can expand the above expression perturbatively to order $g^{2}$ to obtain
@@ C.1.   High-Energy Astrophysics
Making use of the changes of variable $y^{2}\mapsto t+b^{2}$, the integrals above can be expressed in terms of the hypergeometric representation introduced in ˜106 as
@@
To simplify our results, we consider the typical regime $\tfrac{\Lambda_{\mathrm{IR}}}{\omega_{p}}\simeq 1$, $\tfrac{\Lambda_{\mathrm{UV}}}{\omega_{p}}\gg 1$, and $\omega_{\star}\gtrsim\omega_{p}$, which after making use of the limit in ˜107 become
@@ C.2   Axion-QED Reheating
Figure 2: Comparison of the exact integrand (solid lines) with the analytical approximation (dashed lines) for different values of $z$ and $\xi$. Note that both integrands are defined over the domain $y\in[\sqrt{1+z^{-2}},\,\infty)$. The left panel shows the bosonic integrand in ˜152, while the right panel shows the fermionic integrand in ˜162 in the right panels. The proposed approximation reproduces the exact integrand with excellent accuracy across different choices of $z$ and $\xi$ successfully capturing both the near-threshold behaviour and the peak structure. The quality of the fit improves with increasing $z$, precisely in the region that dominates the contribution to the resulting integral.
@@
Using the ansatz ˜152 in ˜151, we can write the integral as
@@
where $\text{erfc}(z)$ is the complementary error function defined in ˜103. We can approximate the function above by studying its asymptotic behaviour relative to the combination $\xi z=1$,
@@
From ˜155, we see that for $\xi|\bm{q}|\lesssim 2m$, the spectral number density remains negligible, as it is heavily suppressed by the exponential factor. This allows us to approximate the total number density as

### v2

@@ 2.1 Vacuum Instability and Pair Production
where the D’Alembertian is $\Box\equiv\partial^{\mu}\partial_{\mu}$. The bosonic functional determinant arises from integrating out charged scalars, while the fermionic one corresponds to integrating out Dirac fields and includes an additional trace over spinor indices. A convenient tool to evaluate these is the proper-time representation (see ˜104). For the scalar case, one has
@@ 3 Stochastic Schwinger Effect
$\Gamma_{\text{1-loop}}^{b}[A_{\mu}]$ $=i\text{Tr}\ln\left[\frac{(\hat{p}-gQ\hat{A}({x}))^{2}+m^{2}-i\epsilon}{\hat{p}^{2}+m^{2}-i\epsilon}\right].$ (26)
@@ 3.1 Stationary background
By inserting ˜124 into ˜31, the decay probability appearing in ˜31 takes the form
@@
where, for shorthand, we denote $\int_{q}\equiv\int d^{4}q$. Here we summarise only the final expressions for the decay into scalar particles; the detailed derivation leading to this result is presented in Appendix B. Using ˜141 and 142, we obtain the one-loop vacuum decay probability into charged bosons induced by a stationary, stochastic gauge field as
@@
where $q^{\mu}=(\omega,\bm{q})$. Using ˜144, the corresponding spectral number density for fermionic pair production takes the form
@@ 3.2 Non-stationary background
with $\langle\!\!\langle\hat{F}_{\mu\nu}(t^{\prime\prime},\bm{q})\,\hat{F}^{\mu\nu}(t^{\prime},-\bm{q})\rangle\!\!\rangle=(q^{2}\eta^{\mu\nu}-q^{\mu}q^{\nu})\langle\!\!\langle\hat{A}_{\mu}(t^{\prime\prime},\bm{q})\,\hat{A}_{\nu}(t^{\prime},-\bm{q})\rangle\!\!\rangle$. The fermionic case follows from ˜144 as
@@ 4.1 Electromagnetic modes in a cold medium
We now compute the vacuum decay rate into a scalar species of mass $m$ in the presence of a cold–plasma background. This quantity admits exact analytical solutions in terms of the Gaussian hypergeometric function $_{2}F_{1}(a,b;c;z)$. The full derivation and closed forms are presented in Appendix C (see ˜152). For phenomenological clarity, here we report only the simplified expression obtained in the representative regime $\Lambda_{\mathrm{IR}}/\omega_{p}\simeq 1$ and $\Lambda_{\mathrm{UV}}/\omega_{p}\gg 1$; the final result reads
@@ 4.2 Dark photon background
Dark photons are hypothetical massive spin-1 bosons from a hidden $U(1)^{\prime}$ sector, kinetically mixed with the Standard Model hypercharge (see Fabbrichesi_2021 ; Caputo:2021eaa for reviews). In astrophysical and cosmological contexts, stochastic dark-photon fields may form with sizeable energy densities, acting as classical sources for charged-particle production and providing a natural setting for the stochastic Schwinger effect Fedderke_2021 ; Cline:2024wja ; Caputo:2025avc . A massive vector field of mass $m_{A^{\prime}}$ propagates three physical degrees of freedom: two transverse and one longitudinal, with dispersion relation
@@
We now evaluate the vacuum decay rate into a scalar species of mass $m$ in the presence of a cold–plasma background. This quantity also admits an exact analytical expression in terms of the Gaussian hypergeometric functions; the detailed derivation and the resulting closed-form expressions can be found in Appendix C (see ˜153).
@@
The charged particles coupled to the dark photon may belong either to a dark sector, consisting of fermionic or bosonic states, or to the Standard Model, in which case they carry a small effective (millicharge) coupling $gQ=\varepsilon e$. In the dark-sector case, the coupling $gQ$ can naturally be of order unity, allowing a stochastic dark-photon background with mass $m_{A^{\prime}}$ to produce pairs with masses up to $m=m_{A^{\prime}}/2$. For SM particles, however, current experimental and astrophysical constraints on $\varepsilon$ impose stringent limits on such interactions. Collider and electroweak precision measurements yield $\varepsilon\lesssim 10^{-2}\,\text{–}\,10^{-3}$ for $m_{A^{\prime}}$ in the GeV range, while at higher masses, $m_{A^{\prime}}\gtrsim 100~\mathrm{GeV}$, these bounds become less restrictive Fabbrichesi_2021 .
@@ 5 Stochastic Schwinger in Axion–Gauge Field Reheating
where the infrared cut-off corresponds to modes residing within the Hubble horizon. In this work, we focus on early-time dynamics with $\mu_{\bm{q}}t\ll 1$, a regime that captures the most relevant phenomenology while avoiding backreaction. Within the Gaussian-window STFT framework in (41), this corresponds to imposing $\sigma\lesssim\mu_{\bm{q}}^{-1}$. As discussed in section˜3.2, in this context the window width is naturally momentum-dependent, scaling with the oscillation timescale of each mode, $\sigma_{\bm{q}}\sim\omega_{\bm{q}}^{-1}$, in close analogy with adaptive wavelet methods. From ˜40 and 41, together with ˜108 and the identification $\sigma\simeq\mu_{\bm{q}}^{-1}$, the gauge mode functions simplify to
@@
As we are working in the small time domain, we can approximate the periodic function as being nearly constant, $c(t,\bm{q})\approx 1$. Using the above, ˜142, ˜143, and in the early-time regime, i.e. $\xi\lvert\bm{q}\rvert t\ll 1$, we find
@@
where $y=\omega/|\mathbf{q}|$, with $e\simeq 2.7$ and the maximal momentum $q_{\text{max}}=m_{\chi}\xi$ set by the instability window in ˜82. The above expression does not admit a closed-form analytic solution. However, we were able to approximate its behaviour using complementary error functions as detailed in appendix˜C of this work. The explicit form of the spectral number density is presented in ˜162. Its contribution remains negligible for small values of $|\bm{q}|$ and becomes relevant for $|{\bm{q}}|\gtrsim 2m/\xi$, where it takes the following form
@@
By performing the momentum integral, we obtain the number density given in ˜163. In the regime of interest, $\xi\gtrsim 1$, it becomes
@@
Finally, we find the background decay rate to bosons as (see ˜165)
@@
In situations where the electromagnetic stochastic component is superimposed on a homogeneous background, we can decompose the gauge potential as
$\hat{A}_{\mu}(x)=\bar{A}_{\mu}(x)+\delta\hat{A}_{\mu}(x),\qquad\langle\delta\hat{A}_{\mu}(x)\rangle_{s}=0,$ (96)
where $\bar{A}_{\mu}(x)$ generates a static electromagnetic field. The background then contributes the usual non-perturbative Schwinger term, while the fluctuations provide an additional stochastic correction set by their two-point correlator. Since the two sectors arise from independent parts of the gauge field, their contributions to the decay exponent add linearly,
$\Gamma_{\text{total}}=\Gamma_{\text{static}}+\Gamma_{\text{stochastic}}.$ (97)
In this regime, the two sectors remain separable at leading order in $\langle\!\!\langle F_{\mu\nu}F^{\mu\nu}\rangle\!\!\rangle$, so the static background does not modify the stochastic contribution. This contrasts with dynamically assisted case with $\langle\!\!\langle\delta F_{\mu\nu}(x)\rangle\!\!\rangle\neq 0$ with
$F_{\mu\nu}F^{\mu\nu}=(\bar{F}_{\mu\nu}+\delta F_{\mu\nu}(x))(\bar{F}^{\mu\nu}+\delta F^{\mu\nu}(x))\supset 2\bar{F}^{\mu\nu}\delta F_{\mu\nu}(x),$ (98)
where such interference between the background and fluctuations can occur.
To put this discussion into perspective, we sharpen the formulation and scope of the stochastic description by distinguishing between deterministic fast oscillations and ensemble stochasticity, which play distinct roles in the Schwinger process.
  • Emergent stochasticity from deterministic fast oscillations: consider a deterministic background field $A_{\mu}(x,t)=\delta A_{\mu}(x,t)$ with rapid spacetime modulations characterized by oscillation scales $(\ell_{\rm osc},\tau_{\rm osc})$. When these satisfy
    $\ell_{\rm osc}\ll\ell_{\rm Sch},\qquad\tau_{\rm osc}\ll\tau_{\rm Sch},$ (99)
    where $(\ell_{\rm Sch}=\frac{mc^{2}}{eE},\tau_{\rm Sch}=\ell_{\rm Sch}/c)$ are the Schwinger tunneling scales, the field averages over many oscillations during a tunneling event. A coarse-graining over intermediate scales then induces an effective stochastic field with covariance
    $G^{\rm eff}_{\mu\nu}(x)=\overline{\delta A_{\mu}(x_{0})\,\delta A_{\nu}(x_{0}+x)},$ (100)
    even though the underlying field is fully deterministic. In this regime the Schwinger process responds only to these coarse-grained statistics, and the stochastic formalism applies without requiring any intrinsic randomness of the original field.
  • Ensemble (macroscopic) stochasticity: Cosmological electromagnetic fields are stochastic in the measure-theoretic sense: the background field is drawn from an ensemble $\{A(x)\}$ with well-defined correlation functions,
    $\langle A_{\mu}(x)\,A_{\nu}(x^{\prime})\rangle_{s}=G_{\mu\nu}(x-x^{\prime}),$ (101)
    even though each realization is slowly varying in space and time. In this situation, the pair-production rate is obtained from the ensemble average of the Schwinger functional,
    $\langle\Gamma\rangle_{s}=\int\mathcal{D}A\,P[A]\,\Gamma[A],$ (102)
    where $P[A]$ is the probability measure characterizing the stochastic field (regardless of $\tau_{\rm osc}$ relative to $\tau_{\rm Sch}$); instead, one performs an average over realizations rather than over rapid microscopic fluctuations. In other words, the Schwinger process is quasi-local in a single realization, but the physically relevant observable in cosmology is the expectation value $\langle\Gamma\rangle_{s}$. A closely related situation occurs in certain astrophysical environments where the electromagnetic field exhibits strong spatial or temporal variability across many coherence domains. If the physical system (or the observer) effectively samples a large number of such statistically independent patches, then the appropriate prediction is again obtained by averaging over the ensemble of coarse-grained field configurations.
In practice, for astrophysical and cosmological electromagnetic fields, it is predominantly the second situation that applies: the fields vary slowly on microscopic scales but are statistically described by an ensemble distribution across many coherence domains, so the physically meaningful prediction is the ensemble-averaged Schwinger rate.
@@ Appendix A Mathematical tools
These identities hold for generic parameters by analytic continuation, except at gamma-function poles. For n=2n=-2, the integral in ˜113 evaluates to
@@ Appendix B Technical Details of Stochastic Effective Action
[…]
@@
$i\Gamma_{\text{1-loop}}^{b}[A_{\mu}]=-\text{Tr}\ln\left(\hat{\mathbb{I}}+i\hat{O}(A)\,\hat{\delta}^{-}\right).$ (127)
Now, we obtain
@@
where to arrive at the last line we have used the identity ˜125.
We now turn to the proof of ˜30. Starting from ˜128 we can expand the above expression perturbatively to order $g^{2}$ to obtain
@@ C.1.   High-Energy Astrophysics
Making use of the changes of variable $y^{2}\mapsto t+b^{2}$, the integrals above can be expressed in terms of the hypergeometric representation introduced in ˜113 as
@@
To simplify our results, we consider the typical regime $\tfrac{\Lambda_{\mathrm{IR}}}{\omega_{p}}\simeq 1$, $\tfrac{\Lambda_{\mathrm{UV}}}{\omega_{p}}\gg 1$, and $\omega_{\star}\gtrsim\omega_{p}$, which after making use of the limit in ˜114 become
@@ C.2   Axion-QED Reheating
Figure 2: Comparison of the exact integrand (solid lines) with the analytical approximation (dashed lines) for different values of $z$ and $\xi$. Note that both integrands are defined over the domain $y\in[\sqrt{1+z^{-2}},\,\infty)$. The left panel shows the bosonic integrand in ˜159, while the right panel shows the fermionic integrand in ˜169 in the right panels. The proposed approximation reproduces the exact integrand with excellent accuracy across different choices of $z$ and $\xi$ successfully capturing both the near-threshold behaviour and the peak structure. The quality of the fit improves with increasing $z$, precisely in the region that dominates the contribution to the resulting integral.
@@
Using the ansatz ˜159 in ˜158, we can write the integral as
@@
where $\text{erfc}(z)$ is the complementary error function defined in ˜110. We can approximate the function above by studying its asymptotic behaviour relative to the combination $\xi z=1$,
@@
From ˜162, we see that for $\xi|\bm{q}|\lesssim 2m$, the spectral number density remains negligible, as it is heavily suppressed by the exponential factor. This allows us to approximate the total number density as